A company will sell Widgets to consumers at a price of $102 per unit. The variable cost to produce Widgets is $50 per unit. The company expects to sell 17,000 Widgets to consumers each year. The fixed costs incurred each year will be $190,000. There is an initial investment to produce the goods of $2,500,000 which will be depreciated straight line over the 13 year life of the investment to a salvage value of $0. The opportunity cost of capital is 5% and the tax rate is 30%.

What is operating cash flow each year?
Answer #1

Sales = 17000*102 =

1734000

Less: Variable cost = 17000*50 = 850000

Less: Fixed cost = 190,000

EBITDA = 694000

Less: Depreciation = 2500000/13 = 192307.69

Profit before tax = 501692.31

Less: Tax = 150507.69

Net Income = 351184.62

Add back depreciation = 192307.69

Operating Cash flow = 543492.31

Hence Annual OCF = $ 543492.31